OpenStudy (anonymous):

verify the identity. cot (x-π/2) = -tan x

OpenStudy (jdoe0001):

$$\bf \large { cot\left(x - \frac{\pi}{2}\right) \implies \cfrac{cos\left(x - \frac{\pi}{2}\right)}{sin\left(x - \frac{\pi}{2}\right)}\\ \implies \cfrac{cos(x)cos\left(\frac{\pi}{2}\right)+sin(x)sin\left(\frac{\pi}{2}\right)} {sin(x)cos\left(\frac{\pi}{2}\right)-cos(x)sin\left(\frac{\pi}{2}\right)} } $$

OpenStudy (jdoe0001):

now, what's \(\bf cos\left(\frac{\pi}{2}\right) \) ? and what's \(\bf sin\left(\frac{\pi}{2}\right) \) ?

OpenStudy (anonymous):

thank you so much! im so sorry for the long pause

OpenStudy (jdoe0001):

yw